OpenStudy (spring98):

medals!!! The Grand Canyon is approximately 2^9 kilometers long. Mariner Valley is a canyon on Mars that is approximately 2^12 kilometers long. About how many times longer is Mariner Valley than the Grand Canyon? A. 2 B. 2^3 C. 2^21 D.2^108

OpenStudy (madhu.mukherjee.946):

B

OpenStudy (michele_laino):

here you have to compute this ratio: \[\Large \frac{{{2^{12}}}}{{{2^9}}} = {2^{12 - 9}} = ...?\]

OpenStudy (spring98):

12^3 right?

OpenStudy (madhu.mukherjee.946):

yeah

OpenStudy (michele_laino):

yes! it is 2^3

OpenStudy (spring98):

thanks @Michele_Laino

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

thanks! @madhu.mukherjee.946 for your help

OpenStudy (madhu.mukherjee.946):

:)@Michele_Laino your welcome